(17/9) + (41/9) =                            .
(17/9) + (41/9) =                            .
(17/9) + (41/9) =                            .

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


(17/9) + (41/9) = 58/9.
Fractions with same denominators are called like fractions.
Sum of two like fractions = (17 + 41)/9

= 58/9


Thanks